2 people emeka and ngozi are 165 metres apart along a line at right angle to a wall which is 165 metres behind ngozi and emeka fires a shoot. What is the 2 report heard by emeka
1
Expert's answer
2020-10-19T13:22:49-0400
"t_1=2\frac{d}{v}=2\frac{165}{330}=1\ s"

"t_1=2\frac{\sqrt{d^2+d^2}}{v}=2\sqrt{2}\frac{165}{330}=1.4\ s"